OpenStudy (anonymous):

Find the derivative of x^(x+17) y ' = _______ how do I start this?

OpenStudy (anonymous):

oh and you simplify it from there to get just (x+17)(x^(x+16)) ?

OpenStudy (anonymous):

ahh okay awesome!! thank you!!

OpenStudy (anonymous):

:/ that's not how the power rule works. The power is not constant.

OpenStudy (anonymous):

let \[y=x ^{x+17},\ln y=\ln x ^{x+17}=\left( x+17 \right)\ln x\] diff.w.r.t. x \[\frac{ y' }{ y }=\left( x+17 \right)*\frac{ 1 }{ x }+\ln x*\left( 1 \right)\] \[y'=y \left\{ \frac{ x+17 }{ x }+\ln x \right\}\] \[\frac{ dy }{ dx }=x ^{x+17}\left\{ \frac{ x+17 }{ x }+\ln x \right\}\]

OpenStudy (anonymous):

ohhh okay i see.. thank you!!

OpenStudy (agent0smith):

Whenever you have an x in exponent, you'll very likely have to use ln on both sides, to differentiate the function.

OpenStudy (anonymous):

ahh okay, awesome!Thanks @agent0smith @surjithayer :)

OpenStudy (anonymous):

yw